# Taming the Tail in Class-Conditional GANs: Knowledge Sharing via Unconditional Training at Lower Resolutions

### v1

@@ Abstract
Despite the extensive research on training generative adversarial networks (GANs) with limited training data, learning to generate images from long-tailed training distributions remains fairly unexplored. In the presence of imbalanced multi-class training data, GANs tend to favor classes with more samples, leading to the generation of low-quality and less diverse samples in tail classes. In this study, we aim to improve the training of class-conditional GANs with long-tailed data. We propose a straightforward yet effective method for knowledge sharing, allowing tail classes to borrow from the rich information from classes with more abundant training data. More concretely, we propose modifications to existing class-conditional GAN architectures to ensure that the lower-resolution layers of the generator are trained entirely unconditionally while reserving class-conditional generation for the higher-resolution layers. Experiments on several long-tail benchmarks and GAN architectures demonstrate a significant improvement over existing methods in both the diversity and fidelity of the generated images. The code is available at https://github.com/khorrams/utlo.
00footnotetext: ${}^{\dagger}$ Correspondence: khorrams@oregonstate.edu
@@ 1 Introduction
Despite the advances, past research on cGANs has primarily focused on learning from balanced data. In real-world scenarios, however, data often follows a power-law distribution, with a few classes dominating most of the training data. This is referred to as the “long-tail (LT) problem[53], where most of the training data come from a few classes (referred to as the “head” of the distribution). In contrast, data from a large number of classes rarely occur (referred to as the “tail” of the distribution). This imbalance in data poses a challenge to the effective learning of tail classes, as standard machine learning algorithms tend to favor densely sampled regions of the input distribution.
@@
To address this challenge, we propose leveraging the abundant information inherent in the head classes, with the aim to learn and infuse knowledge from the head classes into the tail ones, thereby enriching the training distribution for the latter. Nevertheless, explicitly disentangling the class-specific information pertaining to head classes from class-independent information, shared between head and tail classes, poses a significant challenge. We make a key observation that the similarity between head and tail instances is often higher at lower resolutions: discriminative features such as shape or specific texture are usually unveiled at higher resolutions. Conversely, information from the lower resolutions tends to be more class-independent such as background, configuration, or direction of objects (Fig. 6), thus can be shared between the head and tail classes.
@@
  • We propose UTLO, a novel knowledge-sharing framework tailored for training cGANs in the long-tailed setup. UTLO allows infusing information from the head classes to the tail classes via an additional unconditional objective applied to the low-resolution part of the cGAN generator. To the best of our knowledge, this work is the first to demonstrate that not all the layers in a cGAN need to receive class-conditional information, i.e., a cGAN featuring a partially unconditional generator.
@@ 2 Related Work
Conditional Generative Adversarial Networks The initial work on conditional image generation using GANs [7] was presented by [28], in which the class conditions were concatenated to the inputs of the generator and discriminator networks. The AC-GAN [32] proposed the use of auxiliary classification in the discriminator. [2] introduced BigGAN and set a milestone in large-scale and high-resolution conditional image generation. The Authors of StyleGAN [17, 16] first map the input noise and class condition to a latent style space, which is then passed to the multiple layers of the generator for image synthesis. More recently, an extension of the StyleGAN called StyleGAN-XL [40] has become the state-of-the-art in conditional image generation on several datasets, including ImageNet and outperforming more complex and time-consuming approaches such as diffusion models [6].
@@
GAN Regularization under Limited Data Training GANs under limited data is challenging as the discriminator can memorize the training samples, resulting in the collapse of the training or quality degradation of the generated images. Recently, data augmentation techniques have been incorporated into GAN training as a means to mitigate this problem [51, 16, 55]. Additionally, [45] introduced a regularization term to the GAN objective, which tracks the predictions of the discriminator for real and generated images using separate moving averages. [20] employed off-the-shelf vision models in an ensemble of discriminators, demonstrating improved performance in both limited-data and large-scale GAN training.
@@
In the context of few-shot image generation, FastGAN [22] proposed a lightweight architecture with a self-supervised discriminator. [33] introduced cross-domain distance consistency in order to transfer the diversity from a source domain to a target domain. Subsequently, [18] proposed a latent-mixup strategy to smooth the latent space through controlled latent interpolation.
@@
Recently, [55, 41] have observed that class-conditioning can cause mode collapse in the limited data regime. Their proposed work learns from limited but balanced/unlabelled data, which is different from our setup: heavily imbalanced long-tail data. [41] proposed Transitional-cGAN, a training strategy that starts with unconditional GAN objective and injects class-conditional information during a transition period. This approach was found to be effective in mitigating mode-collapse in the limited data regime. Our main idea of separating pathways between the low-resolution and high-resolution is novel w.r.t. [41]: to promote knowledge sharing, we modify cGAN architecture such that the lower-resolution part of the generator is entirely unconditional while only the higher-resolution part is conditional and received class-conditional information.
@@
Long-tail Recognition To address the long-tail recognition problem [53, 47], previous research has primarily focused on techniques such as class re-balancing [3, 37, 27], learning algorithm and model design [13, 14, 56], and information augmentation [23, 44, 46]. More related to our work, re-sampling [54, 8] has been widely used to handle class imbalance problems. This involves balancing the number of samples used during training through over-sampling and/or under-sampling. Over-sampling increases the frequency of tail classes [9, 30] while under-sampling reduces the imbalance by reducing the frequency of head class instances [24, 48]. Note, our research diverges from prior studies on long-tail recognition by addressing the more challenging task of generative modeling of the long-tail data, as opposed to the conventional long-tail classification task.
@@
Training GANs on Long-tail Data Recent years have seen a growing interest in learning from long-tail data. However, most focus has been directed toward recognition tasks, with a limited number of studies addressing the development of generative models for long-tail data. GAMO [30] uses adversarial training to over-sample from minority classes. [34] introduced class-balancing regularization to the GAN objective, which leverages the predictions of a pretrained classifier to improve the generation of underrepresented classes in an unconditional setting. However, this method trains an unconditional GAN and is restricted to having access to a pre-trained classifier on the long-tail data to guide the training. The most closely related work to ours is the Group Spectral Regularization (GSR) [35] and NoisyTwins [36] regularizations. The authors identified mode collapse on tail classes as a result of the spectral norm explosion in class-conditional BatchNorms and the collapse of the latents in the StyleGANs $\mathcal{W}$ space, respectively. In contrast, we present a novel approach for sharing knowledge between head and tail classes to enhance long-tail learning — which is orthogonal to regularization techniques. In addition, while [36] is constrained to StyleGANs, our framework can be extended to different GAN architectures.
@@ 3.1 Background: Conditional Generative Adversarial Networks
Conditional Generative Adversarial Networks (cGANs) are suitable for many applications due to providing user control over the generated samples at inference time. This is particularly helpful in the long-tail setup, where it is desirable to explicitly generate samples from rare (tail) classes. In contrast, unconditional GANs are likely to generate samples that track the training distribution, resulting in most samples coming from common (head) classes. To train a cGAN over a dataset $\mathcal{D}$ containing $n$ instances ${\mathbb{X}}=\{{\bm{x}}_{1},\dots,{\bm{x}}_{n}\}$ and their corresponding labels ${\mathbb{Y}}=\{y_{1},\dots,y_{n}\}$ , we can formulate the adversarial training objective that alternatively optimizes the following loss terms,
@@ 3.2 Caveat of Training cGANs on Long-tail Data
Although regularizations introduced in [35, 36] can help mitigate the collapse of learning in tail classes that contain a sufficiently large number of samples, we still observe mode-collapse occurring when a very limited number of samples are in tail classes — which is inherently challenging. This can be seen in Fig. 2 for tail classes of CIFAR100-LT with as few as 5 training instances. Although early-stopping can be adopted to obtain reasonable performance, the key in boosting the performance would be an approach to infuse additional diversities into the tail classes rather than only relying on techniques such as regularization.
Figure 2: Convergence of the different methods on CIFAR100-LT ($\rho=100$), where tail classes have as few as 5 training examples. Incorporating our framework into the baseline alleviates overfitting as a result of knowledge sharing from head to rare tail classes.
Figure 3: The proposed framework, UTLO, illustrated for the StyleGAN2-ADA architecture. Low and high resolution image paths are used for unconditional and class-conditional objectives, respectively. $z$ is the input latent code, $y$ indicates the class embeddings, and $c$ is a constant input. Separate style vectors, $w_{z}$ (class-independent) and $w_{z,y}$ (class-conditional), are generated using the same z and a shared style-mapping network $G_{\text{map}}$ which are then passed to $G_{l}$ and $G_{h}$, respectively. The high-resolution generated image $\hat{x}\in\mathbb{R}^{3\times H\times H}$ is passed through the discriminator $D=D_{l}\circ D_{h}$ to calculate the conditional objective $\mathcal{L}_{c}$ while the low-resolution image $\hat{x}_{l}\in\mathbb{R}^{3\times L\times L}$ is passed only through $D_{l}$ to calculate the unconditional objective $\mathcal{L}_{uc}$. The final objective $\mathcal{L}$ is the combination of the two. While $D_{l}$ is shared, two separate prediction heads (FC layers) are used for unconditional and conditional objectives. The fromRGB is designed to increase RGB channels to match the input channels of the $D_{l}$.
@@ 3.3 Knowledge Sharing via Unconditional Training at Lower Resolutions (UTLO)
We devise a novel generative adversarial network (GAN) training scheme for long-tail data that aims to share knowledge from predominant head classes to rare tail classes, assuming some of the tail and head classes are at least coarsely similar. We propose a framework that uses unconditional learning at the low-resolution images/features at an intermediate layer of the generator using the unconditional GAN objective. This helps with learning universal features from both head and tail classes. The subsequent layers, responsible for introducing finer details at higher resolutions, are trained using a class-conditional GAN objective.
In our proposed framework, we impose conditional training on the high-resolution image output from the final layer of the generator network. This is built upon the features learned at lower layers through the use of an unconditional objective, inheriting features primarily from head classes. The conditioning on the class labels gives control at the inference time to explicitly generate images from tail classes, meeting our design desideratum. The combination of unconditional and conditional objectives using our proposed method enables knowledge sharing, which in turn improves the training quality of GANs on long-tail data.
@@ 3.3.1 Modifying the Generator
In the generator design of the StyleGAN2-ADA [16, 17], the class-conditional information, combined with the latent vector, is first embedded in the style space $w$ using the style-mapping network $G_{\text{map}}$. The style vectors are then broadcasted across the layers of the synthesis network $G_{\text{syns}}$ in order to generate images. To meet our design desideratum, we dissect the synthesis network into two sub-networks $G_{\text{syns}}=G_{h}\circ G_{l}$ where $G_{l}$ represents the earlier part of the network that produces intermediate features and/or images at a low resolution $L\times L$. $G_{h}$ on the other hand represents the latter part of the $G_{\text{syns}}$ that generates the output image at the ultimate high-resolution $H\times H$ where $H>L$.
To block the flow of class-conditional information to the lower layers of $G_{\text{syns}}$, we generate separate $w$ vectors, one containing the class-conditional embeddings, referred to as $w_{z,y}$, while for the other one, referred to as $w_{z}$, class information was set to zero. Note, both vectors share the same latent variable $z$ and mapping network $G_{\text{map}}$ parameters, and the only difference is the presence of class-conditional information. The lower layers of the synthesis network $G_{l}$ are conditioned on $w_{z}$ while the subsequent layers at higher resolutions receive $w_{z,y}$ (see Fig. 3).
Generators typically follow a network design that gradually increases the resolution of intermediate features and/or images as the network progresses. This allows us to select a desired intermediate low-resolution, such as $8\times 8$ or $16\times 16$, during training. Recent generator designs often incorporate skip connections [15] or residual connections [29] to improve gradient flow. The generator of StyleGAN2-ADA uses skip connections, which explicitly generates intermediate RGB images at each layer. We take this low-resolution image as the input to the discriminator. Note, in the case of a generator design utilizing residual connections, a $1\times 1$ convolutional layer can be incorporated to reduce the number of channels to RGB.
In a forward pass through the generator, images at low and high resolutions can be generated simultaneously with no additional overhead. By inputting a latent code $z\sim p_{z}({\bm{z}})$ and a target label $y$, we can obtain an unconditional low-resolution synthetic image $\hat{{\bm{x}}}_{l}\in\mathbb{R}^{3\times L\times L}$ along with a conditional high-resolution synthetic image $\hat{{\bm{x}}}\in\mathbb{R}^{3\times H\times H}$ at the output. While $\hat{{\bm{x}}}$ is trained to be from class $y$, the former has no class-specific constraints. At inference time, the low-resolution intermediate images can be discarded.
@@ 3.3.2 Modifying the Discriminator
Opposite to the generator design, discriminators gradually reduce the resolution of the intermediate features and/or images as the network progresses. In order to pass images to an intermediate low-resolution layer of the discriminator, we dissect it into two sub-networks. Formally, we define the discriminator as $D=D_{l}\circ D_{h}$ where $D_{h}$ represents the earlier part of the original network that takes the images at training resolution $\mathbb{R}^{3\times H\times H}$. On the other hand, $D_{l}$ represents the later part of the discriminator taking a low-resolution input $\mathbb{R}^{C\times L\times L}$ with $C$ channels.
The default discriminator design in the StyleGAN2-ADA uses residual connections. This does not allow the direct passing of RGB inputs to the intermediate layers. To overcome this, we add a $1\times 1$ convolutional layer, referred to as “fromRGB”, that increases the channels of the RGB images to match $C$, the number of input channels of $D_{l}$.
More concretely, $D_{l}$ is a shared network in both the unconditional low-resolution and class-conditional high-resolution image paths (see Fig. 3). In order to obtain separate unconditional and conditional predictions, we use two different fully-connected (FC) layers at the final layer of $D_{l}$. The output of the $D_{h}$ and the low-resolution image $\hat{{\bm{x}}}_{l}$ are passed through the $D_{l}$ which are then used for calculating the class-conditional loss $\mathcal{L}_{c}$ and unconditional loss $\mathcal{L}_{uc}$, respectively.
Note, when passing real images to $D_{l}$ for the unconditional loss, they are first downsized to the low-resolution $L\times L$ using bi-linear interpolation to be comparable with the ones generated from $G_{l}$. They are then passed through the fromRGB layer and consequently $D_{l}$. For the class-conditional loss, the original images are passed to $D$ without any modification.

### v2

@@ Abstract
Despite extensive research on training generative adversarial networks (GANs) with limited training data, learning to generate images from long-tailed training distributions remains fairly unexplored. In the presence of imbalanced multi-class training data, GANs tend to favor classes with more samples, leading to the generation of low quality and less diverse samples in tail classes. In this study, we aim to improve the training of class-conditional GANs with long-tailed data. We propose a straightforward yet effective method for knowledge sharing, allowing tail classes to borrow from the rich information from classes with more abundant training data. More concretely, we propose modifications to existing class-conditional GAN architectures to ensure that the lower-resolution layers of the generator are trained entirely unconditionally while reserving class-conditional generation for the higher-resolution layers. Experiments on several long-tail benchmarks and GAN architectures demonstrate a significant improvement over existing methods in both the diversity and fidelity of the generated images. The code is available at https://github.com/khorrams/utlo.
@@ 1 Introduction
Despite the advances, past research on cGANs has focused primarily on learning from balanced data. In real-world scenarios, however, data often follows a power-law distribution, with a few classes dominating most of the training data. This is referred to as the “long-tail (LT) problem[55], where most of the training data come from a few classes (referred to as the “head” of the distribution). In contrast, data from a large number of classes rarely occur (referred to as the “tail” of the distribution). This imbalance in data poses a challenge to the effective learning of tail classes, as standard machine learning algorithms tend to favor densely sampled regions of the input distribution.
@@
To address this challenge, we propose to leverage the abundant information in the head classes, with the aim to learn and infuse knowledge from the head classes into the tail ones, thereby enriching the training distribution for the latter. However, it is very difficult to explicitly disentangle the class-specific information pertaining to head classes from class-independent information shared between head and tail classes. We make a key observation that the similarity between head and tail instances is often higher at lower resolutions: discriminative features such as shape or specific texture are usually unveiled at higher resolutions. Conversely, information from the lower resolutions tends to be more class-independent such as background, configuration, or direction of objects (Fig. 6), thus can be shared between the head and tail classes.
@@
  • We propose UTLO, a novel knowledge-sharing framework tailored for training cGANs in the long-tailed setup. UTLO allows infusing information from the head classes to the tail classes via an additional unconditional objective applied to the low-resolution part of the cGAN generator. To the best of our knowledge, this work is the first to demonstrate that not all layers in a cGAN need to receive class-conditional information, i.e., a cGAN featuring a partially unconditional generator.
@@ 2 Related Work
Conditional Generative Adversarial Networks The initial work on conditional image generation using GANs [7] was presented by [29], in which the class conditions were concatenated to the inputs of the generator and discriminator networks. The AC-GAN [33] proposed the use of auxiliary classification in the discriminator. [2] introduced BigGAN and set a milestone in large-scale and high-resolution conditional image generation. The authors of StyleGAN [18, 17] first map the input noise and class condition to a latent style space, which is then passed to the multiple layers of the generator for image synthesis. More recently, an extension of the StyleGAN called StyleGAN-XL [41] has become the state-of-the-art in conditional image generation on several datasets and outperforms more complex and time-consuming approaches such as diffusion models [6].
@@
GAN Regularization under Limited Data Training GANs under limited data is challenging as the discriminator can memorize the training samples, resulting in the collapse of the training or quality degradation of the generated images. Recently, data augmentation and regularization techniques have been incorporated into GAN training as a means to mitigate this problem [53, 17, 57, 15, 58, 46]. Additionally, [47] introduced a regularization term to the GAN objective, which tracks the predictions of the discriminator for real and generated images using separate moving averages. [21] employed off-the-shelf vision models in an ensemble of discriminators, demonstrating improved performance in both limited-data and large-scale GAN training.
@@
In the context of few-shot image generation, FastGAN [23] proposed a lightweight architecture with a self-supervised discriminator. [34] introduced cross-domain distance consistency in order to transfer diversity from a source domain to a target domain. Subsequently, [19] proposed a latent-mixup strategy to smooth the latent space through controlled latent interpolation.
@@
Recently, [57, 42] have observed that class-conditioning can cause mode collapse in the limited data regime. Their proposed work learns from limited but balanced/unlabeled data, which is different from our setup: heavily imbalanced long-tail data. [42] proposed Transitional-cGAN, a training strategy that starts with unconditional GAN objective and injects class-conditional information during a transition period. This approach was found to be effective in mitigating mode-collapse in the limited data regime. Our main idea of separating pathways between low-resolution and high-resolution is novel w.r.t. [42]: to promote knowledge sharing, we modify the cGAN architecture so that the lower-resolution part of the generator is entirely unconditional while only the higher-resolution part is conditional and received class-conditional information.
@@
Long-tail Recognition To address the long-tail recognition problem [55, 49], previous research has primarily focused on techniques such as class re-balancing [3, 38, 28], learning algorithm and model design [13, 14, 59], and information augmentation [24, 45, 48]. More related to our work, re-sampling [56, 8] has been widely used to handle class imbalance problems. This involves balancing the number of samples used during training through over-sampling and/or under-sampling. Oversampling increases the frequency of tail classes [9, 31] while under-sampling reduces the imbalance by reducing the frequency of head class instances [25, 50]. Note that our research diverges from prior studies on long-tail recognition by addressing the more challenging task of generative modeling of the long-tail data, as opposed to the conventional long-tail classification task.
@@
Training GANs on Long-tail Data Recent years have seen a growing interest in learning from long-tail data. However, most focus has been directed toward recognition tasks, with a limited number of studies addressing the development of generative models for long-tail data. GAMO [31] uses adversarial training to over-sample from minority classes. [35] introduced class-balancing regularization to the GAN objective, which leverages the predictions of a pretrained classifier to improve the generation of underrepresented classes in an unconditional setting. However, this method trains an unconditional GAN and is restricted to having access to a pre-trained classifier on the long-tail data to guide the training. The most closely related work to ours is the Group Spectral Regularization (GSR) [36] and NoisyTwins [37] regularizations. The authors identified mode collapse on tail classes as a result of the spectral norm explosion in class-conditional BatchNorms and the collapse of the latents in the StyleGANs $\mathcal{W}$ space, respectively. In contrast, we present a novel approach for sharing knowledge between head and tail classes to enhance long-tail learning, which is orthogonal to regularization techniques. In addition, while [37] is restricted to StyleGANs, our framework can be extended to different GAN architectures.
@@ 3.1 Background: Conditional Generative Adversarial Networks
Conditional Generative Adversarial Networks (cGANs) are suitable for many applications due to providing user control over the generated samples at inference time. This is particularly helpful in the long-tail setup, where it is desirable to explicitly generate samples from rare (tail) classes. In contrast, unconditional GANs are likely to generate samples that track the training distribution, resulting in most samples coming from common (head) classes. To train a cGAN over a dataset containing $n$ instances ${\mathbb{X}}=\{{\bm{x}}_{1},\dots,{\bm{x}}_{n}\}$ and their corresponding labels ${\mathbb{Y}}=\{y_{1},\dots,y_{n}\}$ , we can formulate the adversarial training objective that alternatively optimizes the following loss terms,
@@ 3.2 Caveat of Training cGANs on Long-tail Data
Although regularizations introduced in [36, 37] can help mitigate the collapse of learning in tail classes that contain a sufficiently large number of samples, we still observe mode collapse occurring when a very limited number of samples are in tail classes, which is inherently challenging. This can be seen in Fig. 2 for tail classes of CIFAR100-LT with as few as 5 training instances. Although early-stopping can be adopted to obtain reasonable performance, the key in boosting the performance would be an approach to infuse additional diversities into the tail classes rather than only relying on techniques such as regularization.
Figure 2: Convergence of different methods on CIFAR100-LT ($\rho=100$), where tail classes have as few as 5 training examples. Incorporating our framework into the baseline alleviates overfitting as a result of knowledge sharing from head to rare tail classes.
Figure 3: The proposed framework, UTLO, illustrated for the StyleGAN2-ADA architecture. Low and high resolution image pathways are used for unconditional and class-conditional objectives, respectively. $z$ is the input latent code, $y$ indicates the class embeddings, and $c$ is a constant input. Separate style vectors, $w_{z}$ (class-independent) and $w_{z,y}$ (class-conditional), are generated using the same z and a shared style-mapping network $G_{\text{map}}$ which are then passed to $G_{l}$ and $G_{h}$, respectively. The high-resolution generated image $\hat{x}\in\mathbb{R}^{3\times H\times H}$ is passed through the discriminator $D=D_{l}\circ D_{h}$ to calculate the conditional objective $\mathcal{L}_{c}$ while the low-resolution image $\hat{x}_{l}\in\mathbb{R}^{3\times L\times L}$ is passed only through $D_{l}$ to calculate the unconditional objective $\mathcal{L}_{uc}$. The final objective $\mathcal{L}$ is the combination of the two. While $D_{l}$ is shared, two separate prediction heads (FC layers) are used for unconditional and conditional objectives. The fromRGB is designed to increase the dimensionality of RGB channels to match the input channels of the $D_{l}$.
@@ 3.3 Knowledge Sharing via Unconditional Training at Lower Resolutions (UTLO)
We devise a novel generative adversarial network (GAN) training scheme for long-tail data that aims to share knowledge from predominant head classes to rare tail classes, assuming some of the tail and head classes are at least coarsely similar. We propose a framework that uses unconditional learning in low-resolution images / features at an intermediate layer of the generator using the unconditional GAN objective. This helps with learning universal features from both head and tail classes. The subsequent layers, responsible for introducing finer details at higher resolutions, are trained using a class-conditional GAN objective.
In our proposed framework, we perform conditional training on the high-resolution image output from the final layer of the generator network. This is built upon the features learned at lower layers through the use of an unconditional objective, inheriting features primarily from head classes. The conditioning on the class labels gives control at the inference time to explicitly generate images from tail classes, meeting our design desideratum. The combination of unconditional and conditional objectives using our proposed method enables knowledge sharing, which, in turn, improves the quality of GAN training on long-tail data.
@@ 3.3.1 Modifying the Generator
In the generator design of the StyleGAN2-ADA [17, 18], the class-conditional information, combined with the latent vector, is first embedded in the style space $w$ using the style-mapping network $G_{\text{map}}$. The style vectors are then broadcasted across the layers of the synthesis network $G_{\text{syns}}$ in order to generate diverse images. To meet our design desideratum, we dissect the synthesis network into two sub-networks $G_{\text{syns}}=G_{h}\circ G_{l}$ where $G_{l}$ represents the earlier part of the network that produces intermediate features and/or images at a low resolution $L\times L$. $G_{h}$ on the other hand, represents the latter part of $G_{\text{syns}}$ that generates the output image at the high resolution $H\times H$ where $H>L$.
To block the flow of class-conditional information to the lower layers of $G_{\text{syns}}$, we generate separate $w$ vectors, one containing the class-conditional embeddings, referred to as $w_{z,y}$, while for the other one, referred to as $w_{z}$, class information was set to zero. Note that both vectors share the same latent variable $z$ and mapping network $G_{\text{map}}$ parameters, and the only difference is the presence of class-conditional information. The lower layers of the synthesis network $G_{l}$ are conditioned on $w_{z}$ while the subsequent layers at higher resolutions receive $w_{z,y}$ (see Fig. 3).
Generators typically follow a network design that gradually increases the resolution of intermediate features and/or images as the network progresses. This allows us to select a desired low-resolution, such as $8\times 8$ or $16\times 16$, during training. Recent generator designs often incorporate skip connections [16] or residual connections [30] to improve gradient flow. The generator of StyleGAN2-ADA uses skip connections, which explicitly generates intermediate RGB images at each layer. We take this low-resolution image as the input to the discriminator. Note that in the case of a generator design that uses residual connections, a $1\times 1$ convolutional layer can be incorporated to convert the residual channels to RGB channels.
In a forward pass through the generator, images at low and high resolutions can be generated simultaneously without additional overhead. By inputting a latent code $z\sim p_{z}({\bm{z}})$ and a target label $y$, we can obtain an unconditional synthetic image $\hat{{\bm{x}}}_{l}\in\mathbb{R}^{3\times L\times L}$ along with a conditional high-resolution synthetic image $\hat{{\bm{x}}}\in\mathbb{R}^{3\times H\times H}$ at the output. While $\hat{{\bm{x}}}$ is trained to be from class $y$, $\hat{{\bm{x}}}_{l}$ has no class-specific constraints. At inference time, the intermediate images can be discarded.
@@ 3.3.2 Modifying the Discriminator
Opposite to the generator design, discriminators gradually reduce the resolution of the intermediate features as the network progresses. In order to pass images to an intermediate layer of the discriminator, we dissect it into two sub-networks. Formally, we define the discriminator as $D=D_{l}\circ D_{h}$ where $D_{h}$ represents the earlier part of the original network that takes the images at training resolution $\mathbb{R}^{3\times H\times H}$. On the other hand, $D_{l}$ represents the latter part of the discriminator taking a lower resolution input $\mathbb{R}^{C\times L\times L}$ with $C$ channels.
The default discriminator design in StyleGAN2-ADA uses residual connections. This does not allow direct passing of RGB inputs to the intermediate layers. To overcome this, we add a $1\times 1$ convolutional layer, referred to as “fromRGB”, that increases the channels of the RGB images to match $C$, the number of input channels of $D_{l}$.
More concretely, $D_{l}$ is a shared network in both the unconditional low-resolution and class-conditional high-resolution image pathways (see Fig. 3). To obtain separate unconditional and conditional predictions, we use two different fully-connected (FC) layers as the final layers of $D_{l}$. The output of $D_{h}$ and the low-res image $\hat{{\bm{x}}}_{l}$ are passed through the $D_{l}$ which are then used for calculating the class-conditional loss $\mathcal{L}_{c}$ and the unconditional loss $\mathcal{L}_{uc}$, respectively.
Note that when passing real images to $D_{l}$ for the unconditional loss, they are first downsized to the low-resolution $L\times L$ using bi-linear interpolation to be comparable with the ones generated from $G_{l}$. They are then passed through the fromRGB layer and consequently $D_{l}$. For the class-conditional loss, the original images are passed to $D$ without any modification.